At what point(s) on the circle x
2 + y
2 = 13 is its tangent line parallel
to the line 3x ⠈ ’ 2y =
1
Expert's answer
2021-10-12T10:36:24-0400

Solution:

Let the given circle be "x^2+y^2=13" ...(i)

And line is "3x-2y=6" ...(ii)

Differentiate (i) w.r.t. "x"

"2x+2y.y'=0
\\\Rightarrow x+y.y'=0
\\\Rightarrow y'=-\dfrac xy"

Slope of line in (ii) is "m=\dfrac32"

Since, given line is parallel to the tangent, so their slopes are equal.

"y'=m
\\\Rightarrow -\dfrac xy=\dfrac32
\\\Rightarrow x=-\dfrac 32y\ ...(iii)"

Put (iii) in (i)

"(\dfrac 32y)^2+y^2=13
\\\Rightarrow \dfrac 94y^2+y^2=13
\\\Rightarrow \dfrac{13}4y^2=13
\\\Rightarrow y^2=4
\\\Rightarrow y=\pm2"

Put these values of y in (iii)

When y=2, x=-3

When y=-2, x=3

Thus, points are "(-3,2),(3,-2)".